# Emerging Activity Temporal Hypergraph (EATH), a model for generating realistic time-varying hypergraphs

### v1

@@ VII.2 Burstiness parameter
where $r=\sigma/\langle t\rangle$. Note that this definition takes into account a correction for the number of observed events $n$, so that $B=-1$ for the perfectly regular case and $B=0$ in case of an exponential distribution [21, 58]. Here we consider $\Delta B=B-B_{exp}(t_{m},\langle t\rangle)$, comparing the empirical burstiness $B$ with that of an exponential distribution $P(t)=\theta(t-t_{m})\frac{1}{\langle t\rangle-t_{m}}e^{-\frac{t-t_{m}}{\langle t\rangle-t_{m}}}$ with the same mean value $\langle t\rangle$ and lower cut-off $t_{m}$ (where $\theta(x)$ is the Heaviside function). In particular:
@@ VII.3 EATH groups selection mechanisms
      at the $n$-th step, the node $k$ is chosen with probability $x_{k}=\beta a_{t}(k)\varphi_{k}(m)\prod\limits_{s\in e}\tilde{\omega}_{0}(s,k)\tilde{\omega}_{0}(k,s)$, thus accounting for the long-term memory between $k$ and all the nodes in the incomplete hyperedge $e=\{i,j,...\}$ of size $n-1$, with $\beta=1/\sum_{l\in\mathcal{V}}a_{t}(l)\varphi_{l}(m)\prod\limits_{s\in e}\tilde{\omega}_{0}(s,l)\tilde{\omega}_{0}(l,s)$. Then $k$ is added to $e$.
@@
  • with probability $(1-p)$, the hyperedge is generated starting from a hyperedge already active at time $t-1$. A hyperedge $e^{\prime}$ that exists at time $t-1$ can be the source of the new hyperedge $e$ at time $t$ only if its size $|e^{\prime}|\in[m,m\pm 1]$, i.e. if $e^{\prime}\in\mathcal{E}_{t-1}^{m,m\pm 1}$. In particular, the hyperedge generation $e^{\prime}\to e$ can occur with three different continuation mechanisms, depending on the size of $e^{\prime}$:
@@ VII.4 Extracting EATH parameters from data
The long-term memory matrix $\omega_{0}$ is generated in a transient time through the model dynamics, so that it is compatible with the properties of the nodes, e.g. their activities. We fix an initial weight matrix $w$, where $w(i,j)=(1-\delta_{i,j})$ (with $\delta$ the Kronecker delta) so that initially all the pairs of nodes have equal memory; then we generate the temporal hypergraph with its dynamics by using the weight matrix $w$ as the memory matrix and by updating it at each time-step: the element $w(i,j)$ is simply increased by one every time $i,j$ meet in a hyperedge of any size. The evolution continues until the heterogeneity of the weight matrix reaches $\langle w^{2}\rangle/\langle w\rangle^{2}=g\langle w_{l}^{2}\rangle/\langle w_{l}\rangle^{2}$, where $w_{l}$ is the weight matrix of the weighted projected aggregated network $\mathcal{G}$ of the empirical dataset. Then the frozen memory matrix of the model is set to $\omega_{0}=w$. The parameter $g$ is tuned for each dataset to reproduce the evolution of the weight distribution and its heterogeneity, by setting it inversely proportional to $\eta E$, where $\eta=|\mathcal{L}|/\binom{N}{2}$ is the density in the aggregated projected graph $\mathcal{G}=(\mathcal{V},\mathcal{L})$, and $E=\sum_{t=1}^{T/\delta t}E_{t}$ is the total number of interactions activated. We have indeed observed that this approach reproduces well the empirical evolution of the weights (see SM for more details).
Finally, for each node we define an active sequence of events as a set of consecutive events whose time-distance is lower than $\overline{\Delta}$ and whose sum of the durations is larger than $\overline{\Delta}$: the node is then considered as in the high-activity phase when it is involved in such an active sequence of events, otherwise it is in the low-activity phase. Then we estimate the average time that nodes spend in the high-, $\Delta_{h}$, and low-activity phases, $\Delta_{l}$. $\overline{\Delta}$ depends on the dataset, and is set inversely proportional to $\eta E$ (see SM for more details).
@@ VII.5 Higher-order SIR process simulations
We simulate the higher-order SIR process on a temporal hypergraph =$\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{T/\delta t}$. Initially, we consider a fully susceptible population with only one infected node chosen uniformly at random in the entire population. The epidemic starts at time $t_{0}$ and is simulated until there are no more infected: if the epidemic is not ended when the network reaches its last snapshot, we repeat the network from the initial snapshot, taking into account periodicity due to nights or weekends.
For each of the parameters’ configurations and temporal hypergraphs, we run simulations to estimate the epidemic final size $R_{\infty}$ and the basic reproduction number $R_{0}$. The initial simulation time $t_{0}$ is chosen uniformly at random in the entire time-span $[0,T/\delta t]$. $R_{\infty}$ is obtained by considering only the simulations that produce at least one infection, while $R_{0}$ considering also the simulations in which no transmission occurs [32]. The simulations to observe the temporal dynamics of the epidemic, $I(t)$ and $R(t)$, are instead run using as initial time $t_{0}=0$.

### v2

@@ VII.2 Burstiness parameter
where $r=\sigma/\langle t\rangle$. Note that this definition takes into account a correction for the number of observed events $n$, so that B=1B=-1 for the perfectly regular case and $B=0$ in case of an exponential distribution [21, 58]. Here we consider $\Delta B=B-B_{exp}(t_{m},\langle t\rangle)$, comparing the empirical burstiness $B$ with that of an exponential distribution $P(t)=\theta(t-t_{m})\frac{1}{\langle t\rangle-t_{m}}e^{-\frac{t-t_{m}}{\langle t\rangle-t_{m}}}$ with the same mean value $\langle t\rangle$ and lower cut-off $t_{m}$ (where $\theta(x)$ is the Heaviside function). In particular:
@@ VII.3 EATH groups selection mechanisms
      at the $n$-th step, the node $k$ is chosen with probability $x_{k}=\beta a_{t}(k)\varphi_{k}(m)\prod\limits_{s\in e}\tilde{\omega}_{0}(s,k)\tilde{\omega}_{0}(k,s)$, thus accounting for the long-term memory between $k$ and all the nodes in the incomplete hyperedge $e=\{i,j,$...\} of size n1n-1, with $\beta=1/\sum_{l\in\mathcal{V}}a_{t}(l)\varphi_{l}(m)\prod\limits_{s\in e}\tilde{\omega}_{0}(s,l)\tilde{\omega}_{0}(l,s)$. Then $k$ is added to $e$.
@@
  • with probability (1p)(1-p), the hyperedge is generated starting from a hyperedge already active at time t1t-1. A hyperedge $e^{\prime}$ that exists at time t1t-1 can be the source of the new hyperedge $e$ at time $t$ only if its size $|e^{\prime}|\in[m,m\pm 1]$, i.e. if $e^{\prime}\in\mathcal{E}_{t-1}^{m,m\pm 1}$. In particular, the hyperedge generation $e^{\prime}\to e$ can occur with three different continuation mechanisms, depending on the size of $e^{\prime}$:
@@ VII.4 Extracting EATH parameters from data
The long-term memory matrix $\omega_{0}$ is generated in a transient time through the model dynamics, so that it is compatible with the properties of the nodes, e.g. their activities. We fix an initial weight matrix $w$, where $w(i,j)=(1-\delta_{i,j})$ (with $\delta$ the Kronecker delta) so that initially all the pairs of nodes have equal memory; then we generate the temporal hypergraph with its dynamics by using the weight matrix $w$ as the memory matrix and by updating it at each time-step: the element $w(i,j)$ is simply increased by one every time $i,j$ meet in a hyperedge of any size. The evolution continues until the heterogeneity of the weight matrix reaches $\langle w^{2}\rangle/\langle w\rangle^{2}=g\langle w_{l}^{2}\rangle/\langle w_{l}\rangle^{2}$, where $w_{l}$ is the weight matrix of the weighted projected aggregated network $\mathcal{G}$ of the empirical dataset. Then the frozen memory matrix of the model is set to $\omega_{0}=w$. The parameter $g$ is tuned for each dataset to reproduce the evolution of the weight distribution and its heterogeneity, by setting it inversely proportional to $\eta E$, where $\eta=|\mathcal{L}|/\binom{N}{2}$ is the density in the aggregated projected graph $\mathcal{G}=(\mathcal{V},\mathcal{L})$, and $E=\sum_{t=1}^{\mathcal{T}/\delta t}E_{t}$ is the total number of interactions activated. We have indeed observed that this approach reproduces well the empirical evolution of the weights (see SM, Sec. IA,B, for more details).
Finally, for each node we define an active sequence of events as a set of consecutive events whose time-distance is lower than $\overline{\Delta}$ and whose sum of the durations is larger than $\overline{\Delta}$: the node is then considered as in the high-activity phase when it is involved in such an active sequence of events, otherwise it is in the low-activity phase. Then we estimate the average time that nodes spend in the high-, $\Delta_{h}$, and low-activity phases, $\Delta_{l}$. $\overline{\Delta}$ depends on the dataset, and is set inversely proportional to $\eta E$ (see SM, Sec. IA, for more details). Once obtained $\Delta_{h}$ and $\Delta_{l}$, the coefficients $\varrho_{h}$ and $\varrho_{l}$ which set the nodes transition rates (see Eqs. (2),(3)) can be derived. In particular $\varrho_{h}$ and $\varrho_{l}$ can be obtained, by inverting Eqs. (2),(3) and considering the definitions of $\Delta_{l}=\langle\frac{1}{r_{l\to h}(i,t)}\rangle$ and of $\Delta_{h}=\langle\frac{1}{r_{h\to l}(i,t)}\rangle$, as:
$\varrho_{l}=\frac{1}{\Delta_{l}}\left\langle\frac{1}{\Lambda_{t}}\right\rangle\left\langle\frac{1}{a_{T}}\right\rangle\langle a_{T}\rangle\ ,\varrho_{h}=\frac{1}{\Delta_{h}}\left\langle\frac{1}{1-\Lambda_{t}}\right\rangle,$
where the average is either over nodes (for $a_{T}$ and $1/a_{T}$) or over time (for $1/\Lambda_{t}$ and $1/(1-\Lambda_{t})$).
@@ VII.5 Higher-order SIR process simulations
We simulate the higher-order SIR process on a temporal hypergraph $\mathscr{H}=\{\mathcal{H}_{t}\}_{t=1}^{\mathcal{T}/\delta t}$. Initially, we consider a fully susceptible population with only one infected node chosen uniformly at random in the entire population. The epidemic starts at time $t_{0}$ and is simulated until there are no more infected: if the epidemic is not ended when the network reaches its last snapshot, we repeat the network from the initial snapshot, taking into account periodicity due to nights or weekends.
For each of the parameters’ configurations and temporal hypergraphs, we run simulations to estimate the epidemic final size $R_{\infty}$ and the basic reproduction number $R_{0}$. The initial simulation time $t_{0}$ is chosen uniformly at random in the entire time-span $0,\mathcal{T}]$. $R_{\infty}$ is obtained by considering only the simulations that produce at least one infection, while $R_{0}$ considering also the simulations in which no transmission occurs [32]. The simulations to observe the temporal dynamics of the epidemic, $I(t)$ and $R(t)$, are instead run using as initial time $t_{0}=0$.